dy/dx + y/x =y-2


dy/dx +xy=x sqrt y



1
Expert's answer
2021-11-04T17:13:35-0400

a. Rewrite in the form of a first order Bernoulli ODE


"\dfrac{dy}{dx}+\dfrac{1}{x}y=y^{-2}"

Substitution "z=y^{1-(-2)}=y^3"


"\dfrac{dz}{dx}=3y^2\dfrac{dy}{dx}"

"y^2\dfrac{dy}{dx}+\dfrac{1}{x}y^3=1"

"\dfrac{1}{3}\dfrac{dz}{dx}+\dfrac{1}{x}z=1"

"\dfrac{dz}{dx}+\dfrac{3}{x}z=3"

Integrating factor


"\mu(x)=e^{\int(3/x)dx}=x^3"

"x^3\dfrac{dz}{dx}+3x^2z=3x^3"

"d(x^3z)=3x^3dx"

Integrate


"\int d(x^3z)=\int 3x^3dx"

"x^3z=\dfrac{3}{4}x^4+C"

"z=\dfrac{3}{4}x+\dfrac{C}{x^3}"

"y=\sqrt[3]{\dfrac{3}{4}x+\dfrac{C}{x^3}}"


b. The equation is in the form of a first order Bernoulli ODE


"\dfrac{dy}{dx}+xy=xy^{1/2}"

Substitution "z=y^{1-(1\/2)}=y^{1\/2}"


"\dfrac{dz}{dx}=\dfrac{1}{2\sqrt{y}}\dfrac{dy}{dx}"

"\dfrac{1}{2\sqrt{y}}\dfrac{dy}{dx}+\dfrac{1}{2}x\sqrt{y}=\dfrac{1}{2}x"

"\dfrac{dz}{dx}+\dfrac{1}{2}xz=\dfrac{1}{2}x"

Integrating factor


"\mu(x)=e^{\int(x/2)dx}=e^{x^2/4}"

"e^{x^2/4}\dfrac{dz}{dx}+\dfrac{1}{2}xe^{x^2/4}z=\dfrac{1}{2}xe^{x^2/4}"

"d(e^{x^2/4}z)=\dfrac{1}{2}xe^{x^2/4}dx"

Integrate


"\int d(e^{x^2/4}z)=\int \dfrac{1}{2}xe^{x^2/4}dx"

"e^{x^2\/4}z=e^{x^2\/4}+C"

"z=1+Ce^{-x^2\/4}"

"y=(1+Ce^{-x^2\/4})^2"